# citations.bib

### v1

@@ Abstract
This paper presents a novel approach for structure-preserving planar simplification of indoor scene point clouds for both simulated and real-world environments. Initially, the scene point cloud undergoes preprocessing steps, including noise reduction and Manhattan world alignment, to ensure robustness and coherence in subsequent analyses. We segment each captured scene into structured (walls-ceiling-floor) and non-structured (indoor objects) scenes. Leveraging a RANSAC algorithm, we extract primitive planes from the input point cloud, facilitating the segmentation and simplification of the structured scene. The best-fitting wall meshes are then generated from the primitives, followed by adjacent mesh merging with the vertex-translation algorithm which preserves the mesh layout. To accurately represent ceilings and floors, we employ the mesh clipping algorithm which clips the ceiling and floor meshes with respect to wall normals. In the case of indoor scenes, we apply a surface reconstruction technique to enhance the fidelity. This paper focuses on the intricate steps of the proposed scene simplification methodology, addressing complex scenarios such as multi-story and slanted walls and ceilings. We also conduct qualitative and quantitative performance comparisons against popular surface reconstruction, shape approximation, and floorplan generation approaches.
@@ 1 Introduction
Our approach begins with a raw point cloud as input, which undergoes axis alignment, scene segmentation, primitives extraction, and simplification to ultimately generate a polygonal mesh. To simplify the complexity of the scene, we segment the point cloud into structured and non-structured scenes where a structured scene represents floor, walls, and ceiling while non-structured scene refers to indoor objects like tables, desks, beds, and so on. This allows for separate processing of structured and non-structured scenes: structured scenes are simplified through planar primitive extraction and non-structured scenes require surface reconstruction due to intricacies that cannot be represented with planar primitives. We use RANSAC for extracting plane primitives similar to [rs14174275, li2016manhattan, ksr, nan2017polyfit]. However, instead of using candidate faces, hypothesis, and selection strategy [nan2017polyfit, ksr], we select adjacent planes and enclose them towards their intersection. This significantly reduces the computational time and the number of faces which we study in detail in section 5. To enhance the geometric representation of the structured scenes, we introduce two novel algorithms: neighboring mesh vertices translation and mesh clipping. The vertex translation algorithm ensures the enclosedness of neighboring wall meshes by translating their vertices to a common intersection point. To avoid one-to-all mapping between the wall meshes, we maintain an adjacency list for each wall mesh. Mesh clipping algorithm, on the other hand, preserves the geometrical structure of ceiling and floor planes by clipping them with respect to adjacent wall meshes. We assess the performance of our pipeline across simulated, RGBD, and real-world scenes, as described in section  5. We further provide qualitative and quantitative comparisons with popular shape approximation, surface reconstruction, and floorplan estimation approaches.
@@ 2 Related Works
This section provides a comprehensive review of related research on the simplification of 3D indoor environments, including surface reconstruction, shape approximation and floorplan generation.
@@ 2.1 Primitive Extraction and Scene Segmentation
Once primitive planes are detected, segmentation of structured and non-structured scenes involves considering the angles between the normal plane or its points and the principal coordinate axes [wang2018plane, sanchez2012planar]. Geometric and structural constraints are also utilized to identify structural components [rs13193844, fang2021structure]. Modern deep learning approaches such as PointNet++ [qi2017pointnet++], and LFCG-Net [rs14194820] leverage semantic information for indoor scene segmentation without need for primitive extraction.
@@ 2.2 Surface Reconstruction Approaches
Several algorithms focus on effectively partitioning space into polygonal faces, convex polyhedra, or computing tetrahedralizations based on intersections of planar primitives. [nan2017polyfit] proposes a framework for reconstructing lightweight polygonal faces from point clouds based on a hypothesis plane generation followed by selection through linear solvers such as SCIP [BolusaniEtal2024ZR], GLPK [glpk], and Gurobi [gurobi]. However, as the complexity of the scene increases, it struggles to generate an accurate scene representation due to a large number of candidate faces. [ksr] employs a shape assembling mechanism utilizing kinetic data structures for space partitioning into convex polyhedra while Constrained Delaunay Tetrahedralization (CDT) is computed through the intersections of planar primitives [8613683, mura2014reconstructing]. Despite largely mitigating the computational inefficiencies, it still struggles with complex and real-world scenes, compromising the scene geometry and generating a large number of polyhedral faces.
[fang2021structure] adresses this issue by utilizing a 3D partitioning data structure with a global and local slicing strategy based on a three-level hierarchy: extracting ceiling and floor primitives, selecting wall planes, and recovering all small planes adjacent to the walls. This approach generates a compact structure mesh with reconstruction of indoor and outdoor environments. While some algorithms specialize in constructing outdoor buildings [chen2022reconstructing, nan2017polyfit, bouzas2020structure], others focus on indoor properties, leveraging methods like merging boxes based on wall classification [murali2017indoor], Markov Random Field (MRF) formulation [li2016manhattan], or voxel-based occupancy grid [yang2022polygonal]. Additionally, the reconstruction of indoor objects is approached either through traditional surface reconstruction algorithms [fang2021structure, sanchez2012planar] or by replacing objects with existing models using registration [rs14194820].
@@ 2.3 Shape Approximation Approaches
Another approach to generate a simplified mesh is to approximate the shape of indoor environments with a small number of faces based on the original shape. With a dense raw triangular mesh as input, generally created from surface reconstruction algorithm such as Poisson Surface Reconstruction [kazhdan2013poisson], a coarser mesh is approximated by simplifying the faces until a user-defined criterion is satisfied. The criterion can either be a target number of faces or a geometric error between the input and simplified mesh [Garland2023SurfaceSU, 745314]. These approaches depend on the edge-collapse operator to iteratively merge the adjacent faces. Based on this, adaptive thresholding can be further added to emphasize planar surfaces [garland1999quadric]. This quadric error metric (QEM) is enhanced by refining the simplification process to operate planar clusters, effectively preserving plane shapes and sharp features while maintaining mesh integrity [8491005]. Similarly, [rossignac1993vertex] introduces an approach that groups vertices into clusters, and for each cluster, computes a new representative vertex for decimation. The method proposed in [cohenVSA2004] iteratively computes the regions that best fit the corresponding planar parts and adjusts proxies in each region. Constrained Delaunay triangulation [huang20173dlite] is also used to triangulate the planar primitives for accurately preserving the boundaries. However, these shape approximation approaches not being structure-aware may not preserve the original scene geometry.
@@ 2.4 Floorplan Generation Approaches
Floorplan approaches focus only on the generation of a 2D layout. Thus, the simplest way to generate a structured 3D model (also popular as 2.5D model) is to project the 2D layout with a predefined camera height. These approaches generally rely on either scene segmentation from images [lin2019floorplan, 360dfpe, liu2018floornet, chen2019floorsp] (leveraging geometric segmentation [sun2019horizonnet] and semantic segmentation [lee2017roomnet, hirzer2020smart]) or planar primitive patches merging into a planar graph [qi2017pointnet, qi2017pointnet++, he2017mask]. Approaches like [mura2014automatic] construct a cell complex in the 2D floor plane followed by individual room partitioning to generate a room polyhedra.
Our method focuses on the piecewise planar primitive reconstruction for structured scenes along with surface reconstruction for non-structured scenes. Unlike approaches such as [nan2017polyfit, ksr], we avoid the candidate faces/polyhedra generation and linear solvers that significantly reduce the computational complexity and the number of faces. Instead of partitioning 3D space in general [mura2016piecewise, ksr], we consider the nature of planar primitives maintaining the local adjacency leveraged by novel vertex translation and mesh clipping algorithms which further refine the geometry. Inspired by the hierarchy approach of [fang2021structure], we use a two-level hierarchy for scene segmentaion: extracting structured planes into i) floors and ceiling, and ii) walls. However, [fang2021structure] fails to incorporate the slanted ceilings and walls, which we address by generating oriented planar meshes with respect to the Z-axis. In general, our approach reduces the computational costs while still addressing complex scenes like multi-story scenes, and scenes with slanted roofs and walls.
@@ 3 Overview
The aligned point cloud undergoes segmentation into structured and non-structured primitives, which are then processed separately. For structured primitives, due to inaccuracies during data capture and registration, we first generate plane models that best fit the primitive point clouds followed by the generation of a simplified mesh best-representing the plane models. These simplified meshes may not accurately represent the geometry of the structured scene, so we further refine these meshes using vertex translation and mesh clipping algorithms.
@@ 4.1.1 Alignment
For easier scene segmentation and planar mesh estimation, we align the point cloud with respect to the Z-axis (upward direction aligns with the Z-axis) and XY-plane adapted from [li2016manhattan] in accordance with the Manhattan world assumptions [manhattanworld]. This step serves primarily in the planar mesh estimation. Although the structured scenes align as per the Manhattan world assumptions, they may contain the primitives unaligned with the primary axes, such as slanted ceilings and walls. In such cases, we generate oriented meshes directly from the primitives. This provides flexibility in the generation of both the axis-aligned and the oriented meshes. The steps adopted to align a point cloud along the Z-axis are described in algorithm 1. This algorithm derives a rotation matrix $\mathbf{R}$, which, upon application, rotates the point cloud for alignment with the Z-axis.
@@ 4.2.1 Vertex Translation
For $M_{P_{i}}\;\epsilon\;M_{walls}$, this algorithm first determines the adjancent wall meshes from $\mathcal{G}_{i}$. Since the adjacent walls generally intersect with each other, we determine the point of intersection, X = $(x,y,z)$ between $M_{P_{i}}$ plane, its adjancent $M_{P_{j}}$ plane and an aribtrary plane along $Z=0$ as shown in figure 4. Let $\nu_{i}$, $\nu_{j}$ be the plane parameters of $M_{P_{i}}$, $M_{P_{i}}$ with normals $\hat{n}_{i}$, $\hat{n}_{j}$ respectively. Then,
@@
Figure 4: Vertex translation. The vertex $\mathbf{V}_{1}$ nearest to the line of intersection between $P_{i}$ and $P_{j}$ is translated to point $\mathbf{Q}_{1}$. The same process is repeated for vertex $\mathbf{V}_{2}$. $Z=0$ is an aribtrary plane to determine the point of intersection X.
@@ 5.1 Dataset Description
We evaluate our algorithm across a diverse range of datasets, including simulated scenes from the Replica dataset[replica19arxiv], RGBD scenes from HM3D [ramakrishnan2021hm3d] and MP3D [Matterport3D], as well as custom dataset captured using LiDAR from iPad Pro (M1) and processed with RTAB-Map [Labb__2018]. Large scene point clouds for custom dataset are generated using Generalized Iterative-Closest Point (ICP) registration [GeneralizedICP]. In cases of public datasets with scene representations as mesh, we convert the mesh into point cloud, maintaining the original number of vertices by employing poisson-disk sampling [6143943]. We compare our approach against established surface reconstruction, shape approximation, and floorplan generation approaches.
@@ 5.2 Pipeline Evaluation
We comprehensively assess the performance of our approach across a variety of scene complexities, encompassing both simulated and custom dataset scenes. Ranging from simple single-room environments to intricate multi-story scenes, the evaluation demonstrates the adaptability and robustness of our approach. Across all evaluated scenes, our approach consistently generates structure-preserving planar simplified scenes as shown in figures 6, 7, 8 and 9. For mulit-story scenes, we extend our analysis to intermediate floors and ceilings as discussed in Section 4.1.2. However, we acknowledge the suboptimal nature of this method for intermediate planes and we plan to improve this in our future enhancements.
@@
Figure 7: Evaluation on multiple floor scenes from Replica (Simulated) and HM3D (Real-world) datasets. The extended mesh within the black region is due to the presence of non-structred object instead of wall on the corresponding side point clouds.
We further evaluate the performance on custom dataset. Due to the inherent noise from multiple overlapping regions in ICP registration, multiple coplanar planes may be generated for a single planar primitive. We mitigate this challenge through appropriate selection of hyperparameters during primitive extraction. Nonetheless, sparse point clouds can affect the generation of structured planes as shown in figure 8.
Figure 8: Evaluation on custom dataset for different scene complexities. The highlighted regions represent sparse or complete lack of points in the input point clouds. The black rectangular regions represent the absence of meshes due to sparse or absence of point clouds in raw input.
@@ 5.3 Comparison with Surface Reconstruction Approaches
Table 1: Comparison of number of faces, $\mathcal{F}$ (for structured scenes) against surface reconstruction approaches. For all scences, our approach generates significantly lower number of faces without compromising the scene accuracy.
@@
Figure 10: Comparison with surface reconstruction approaches including both structred and non-structred scenes.
@@ 5.4 Comparison with Shape Approximation Approaches
Figure 11: Comparison with shape approximation approaches. The blue region shows the unclipped mesh due to the absence of ceiling in the original mesh.
As shown in figure 11, all approaches generate structure-preserving mesh. However, focusing on structured scenes, the primary objective of our work, we observe that our approach significantly reduces the number of faces, as detailed in table 3, thus largely simplifying the structured mesh. Unlike other approaches, our method inherently distinguishes between structured and non-structured scenes. Thus, for a fair comparison, table 3 contains a comparison of number of faces for only the structured scenes. Notably, in scenes such as frl_apartment_0 in figure 11, where ceilings are absent, our approach’s dependence on clipping ceiling planes with respect to neighboring planes may result in unclipped extended planes, as indicated by the blue region.
Additionally, we evaluate RMSE, as shown in table 4, including both structured and non-structured scenes. RMSE values across all approaches are markedly lower compared to surface reconstruction, owing to the inherent nature of shape approximation approaches, which focus on mesh decimation rather than reconstruction from the original mesh. In simpler scenes like office_4, the lower number of faces in our approach result in lower RMSE. However, as scene complexity increases such as hotel_0, the number of primitives and faces also increases, leading to elevated RMSE values.

### v2

@@ Abstract
This paper presents a novel approach for structure-preserving planar simplification of indoor scene point clouds for both simulated and real-world environments. The scene point cloud initially undergoes preprocessing steps, including noise reduction and Manhattan world alignment, to ensure robustness and coherence in subsequent analyses. We segment each captured scene into structured (walls-ceiling-floor) and non-structured (indoor objects) scenes. Leveraging a RANSAC algorithm, we extract primitive planes from the input point cloud, facilitating the segmentation and simplification of the structured scene. The best-fitting wall meshes are then generated from the primitives, followed by adjacent mesh merging with the vertex-translation algorithm which preserves the mesh layout. To accurately represent ceilings and floors, we employ the mesh clipping algorithm which clips the ceiling and floor meshes with respect to wall normals. In the case of indoor scenes, we apply a surface reconstruction technique to enhance the fidelity. This paper focuses on the intricate steps of the proposed scene simplification methodology, addressing complex scenarios such as multi-story and slanted walls and ceilings. We also compare qualitative and quantitative performance against popular surface reconstruction, shape approximation, and floorplan generation approaches.
@@ 1 Introduction
Our approach begins with a raw point cloud as input, which undergoes axis alignment, scene segmentation, primitives extraction, and simplification to ultimately generate a polygonal mesh. To simplify the complexity of the scene, we segment the point cloud into structured and non-structured scenes where a structured scene represents the floor, walls, and ceiling while non-structured scene refers to indoor objects like tables, desks, beds, and so on. This allows for separate processing of structured and non-structured scenes: structured scenes are simplified through planar primitive extraction and non-structured scenes require surface reconstruction due to intricacies that cannot be represented with planar primitives. We use RANSAC for extracting plane primitives similar to [rs14174275, li2016manhattan, ksr, nan2017polyfit]. However, instead of using candidate faces, hypothesis, and selection strategy [nan2017polyfit, ksr], we select adjacent planes and enclose them towards their intersection. This significantly reduces the computational time and the number of faces which we study in detail in section 5. To enhance the geometric representation of the structured scenes, we introduce two novel algorithms: neighboring mesh vertices translation and mesh clipping. The vertex translation algorithm ensures the enclosedness of neighboring wall meshes by translating their vertices to a common intersection point. To avoid one-to-all mapping between the wall meshes, we maintain an adjacency list for each wall mesh. The mesh clipping algorithm, on the other hand, preserves the geometrical structure of ceiling and floor planes by clipping them with respect to adjacent wall meshes. We assess the performance of our pipeline across simulated, RGBD, and real-world scenes, as described in section  5. We further provide qualitative and quantitative comparisons with popular shape approximation, surface reconstruction, and floorplan estimation approaches.
@@ 2 Related Works
This section provides a comprehensive review of related research on the simplification of 3D indoor environments, including surface reconstruction, shape approximation, and floorplan generation.
@@ 2.1 Primitive Extraction and Scene Segmentation
Once primitive planes are detected, segmentation of structured and non-structured scenes involves considering the angles between the normal plane or its points and the principal coordinate axes [wang2018plane, sanchez2012planar]. Geometric and structural constraints are also utilized to identify structural components [rs13193844, fang2021structure]. Modern deep learning approaches such as PointNet++ [qi2017pointnet++], and LFCG-Net [rs14194820] leverage semantic information for indoor scene segmentation without the need for primitive extraction.
@@ 2.2 Surface Reconstruction Approaches
Several algorithms focus on effectively partitioning space into polygonal faces, convex polyhedra, or computing tetrahedralizations based on intersections of planar primitives. [nan2017polyfit] proposes a framework for reconstructing lightweight polygonal faces from point clouds based on a hypothesis plane generation followed by selection through linear solvers such as SCIP [BolusaniEtal2024ZR], GLPK [glpk], and Gurobi [gurobi]. However, as the complexity of the scene increases, it struggles to generate an accurate scene representation due to the large number of candidate faces. [ksr] employs a shape assembling mechanism utilizing kinetic data structures for space partitioning into convex polyhedra while Constrained Delaunay Tetrahedralization (CDT) is computed through the intersections of planar primitives [8613683, mura2014reconstructing]. Despite largely mitigating the computational inefficiencies, it still struggles with complex and real-world scenes, compromising the scene geometry and generating a large number of polyhedral faces.
[fang2021structure] addresses this issue by utilizing a 3D partitioning data structure with a global and local slicing strategy based on a three-level hierarchy: extracting ceiling and floor primitives, selecting wall planes, and recovering all small planes adjacent to the walls. This approach generates a compact structure mesh with the reconstruction of indoor and outdoor environments. While some algorithms specialize in constructing outdoor buildings [chen2022reconstructing, nan2017polyfit, bouzas2020structure], others focus on indoor properties, leveraging methods like merging boxes based on wall classification [murali2017indoor], Markov Random Field (MRF) formulation [li2016manhattan], or voxel-based occupancy grid [yang2022polygonal]. Additionally, the reconstruction of indoor objects is approached either through traditional surface reconstruction algorithms [fang2021structure, sanchez2012planar] or by replacing objects with existing models using registration [rs14194820].
@@ 2.3 Shape Approximation Approaches
Another approach to generating a simplified mesh is to approximate the shape of indoor environments with a small number of faces based on the original shape. With a dense raw triangular mesh as input, generally created from surface reconstruction algorithm such as Poisson Surface Reconstruction [kazhdan2013poisson], a coarser mesh is approximated by simplifying the faces until a user-defined criterion is satisfied. The criterion can either be a target number of faces or a geometric error between the input and simplified mesh [Garland2023SurfaceSU, 745314]. These approaches depend on the edge-collapse operator to iteratively merge the adjacent faces. Based on this, adaptive thresholding can be further added to emphasize planar surfaces [garland1999quadric]. This quadric error metric (QEM) is enhanced by refining the simplification process to operate planar clusters, effectively preserving plane shapes and sharp features while maintaining mesh integrity [8491005]. Similarly, [rossignac1993vertex] introduces an approach that groups vertices into clusters, and for each cluster, computes a new representative vertex for decimation. The method proposed in [cohenVSA2004] iteratively computes the regions that best fit the corresponding planar parts and adjusts proxies in each region. Constrained Delaunay triangulation [huang20173dlite] is also used to triangulate the planar primitives for accurately preserving the boundaries. However, these shape approximation approaches not being structure-aware may not preserve the original scene geometry.
@@ 2.4 Floorplan Generation Approaches
Floorplan approaches focus only on the generation of a 2D layout. Thus, the simplest way to generate a structured 3D model (also popular as a 2.5D model) is to project the 2D layout with a predefined camera height. These approaches generally rely on either scene segmentation from images [lin2019floorplan, 360dfpe, liu2018floornet, chen2019floorsp] (leveraging geometric segmentation [sun2019horizonnet] and semantic segmentation [lee2017roomnet, hirzer2020smart]) or planar primitive patches merging into a planar graph [qi2017pointnet, qi2017pointnet++, he2017mask]. Approaches like [mura2014automatic] construct a cell complex in the 2D floor plane followed by individual room partitioning to generate a room polyhedra.
Our method focuses on the piecewise planar primitive reconstruction for structured scenes along with surface reconstruction for non-structured scenes. Unlike approaches such as [nan2017polyfit, ksr], we avoid the candidate faces/polyhedra generation and linear solvers that significantly reduce the computational complexity and the number of faces. Instead of partitioning 3D space, in general, [mura2016piecewise, ksr], we consider the nature of planar primitives maintaining the local adjacency leveraged by novel vertex translation and mesh clipping algorithms which further refine the geometry. Inspired by the hierarchy approach of [fang2021structure], we use a two-level hierarchy for scene segmentation: extracting structured planes into i) floors and ceiling, and ii) walls. However, [fang2021structure] fails to incorporate the slanted ceilings and walls, which we address by generating oriented planar meshes with respect to the Z-axis. In general, our approach reduces the computational costs while still addressing complex scenes like multi-story scenes and scenes with slanted roofs and walls.
@@ 3 Overview
The aligned point cloud undergoes segmentation into structured and non-structured primitives, which are then processed separately. For structured primitives, due to inaccuracies during data capture and registration, we first generate plane models that best fit the primitive point clouds followed by the generation of a simplified mesh best representing the plane models. These simplified meshes may not accurately represent the geometry of the structured scene, so we further refine these meshes using vertex translation and mesh clipping algorithms.
@@ 4.1.1 Alignment
For easier scene segmentation and planar mesh estimation, we align the point cloud with respect to the Z-axis (upward direction aligns with the Z-axis) and XY-plane adapted from [li2016manhattan] following the Manhattan world assumptions [manhattanworld]. This step serves primarily in the planar mesh estimation. Although the structured scenes align as per the Manhattan world assumptions, they may contain the primitives unaligned with the primary axes, such as slanted ceilings and walls. In such cases, we generate oriented meshes directly from the primitives. This provides flexibility in the generation of both the axis-aligned and the oriented meshes. The steps adopted to align a point cloud along the Z-axis are described in algorithm 1. This algorithm derives a rotation matrix $\mathbf{R}$, which, upon application, rotates the point cloud for alignment with the Z-axis.
@@ 4.2.1 Vertex Translation
For $M_{P_{i}}\;\epsilon\;M_{walls}$, this algorithm first determines the adjacent wall meshes from $\mathcal{G}_{i}$. Since the adjacent walls generally intersect with each other, we determine the point of intersection, X = $(x,y,z)$ between $M_{P_{i}}$ plane, its adjacent $M_{P_{j}}$ plane and an arbitrary plane along $Z=0$ as shown in figure 4. Let $\nu_{i}$, $\nu_{j}$ be the plane parameters of $M_{P_{i}}$, $M_{P_{i}}$ with normals $\hat{n}_{i}$, $\hat{n}_{j}$ respectively. Then,
@@
Figure 4: Vertex translation. The vertex $\mathbf{V}_{1}$ nearest to the line of intersection between $P_{i}$ and $P_{j}$ is translated to point $\mathbf{Q}_{1}$. The same process is repeated for vertex $\mathbf{V}_{2}$. $Z=0$ is an arbitrary plane to determine the point of intersection X.
@@ 5.1 Dataset Description
We evaluate our algorithm across a diverse range of datasets, including simulated scenes from the Replica dataset[replica19arxiv], RGBD scenes from HM3D [ramakrishnan2021hm3d] and MP3D [Matterport3D], as well as custom datasets captured using LiDAR from iPad Pro (M1) and processed with RTAB-Map [Labb__2018]. Large scene point clouds for custom datasets are generated using Generalized Iterative-Closest Point (ICP) registration [GeneralizedICP]. In the case of public datasets with scene representations as mesh, we convert the mesh into a point cloud, maintaining the original number of vertices by employing poisson-disk sampling [6143943]. We compare our approach against established surface reconstruction, shape approximation, and floorplan generation approaches.
@@ 5.2 Pipeline Evaluation
We comprehensively assess the performance of our approach across a variety of scene complexities, encompassing both simulated and custom dataset scenes. Ranging from simple single-room environments to intricate multi-story scenes, the evaluation demonstrates the adaptability and robustness of our approach. Across all evaluated scenes, our approach consistently generates structure-preserving planar simplified scenes as shown in figures 6, 7, 8 and 9. For multi-story scenes, we extend our analysis to intermediate floors and ceilings as discussed in Section 4.1.2. However, we acknowledge the suboptimal nature of this method for intermediate planes and we plan to improve this in our future enhancements.
@@
Figure 7: Evaluation on multiple floor scenes from Replica (Simulated) and HM3D (Real-world) datasets. The extended mesh within the black region is due to the presence of non-structured objects instead of a wall on the corresponding side point clouds.
We further evaluate the performance of the custom dataset. Due to the inherent noise from multiple overlapping regions in ICP registration, multiple coplanar planes may be generated for a single planar primitive. We mitigate this challenge through the appropriate selection of hyperparameters during primitive extraction. Nonetheless, sparse point clouds can affect the generation of structured planes as shown in figure 8.
Figure 8: Evaluation on the custom dataset for different scene complexities. The highlighted regions represent sparse or complete lack of points in the input point clouds. The black rectangular regions represent the absence of meshes due to sparse or absence of point clouds in raw input.
@@ 5.3 Comparison with Surface Reconstruction Approaches
Table 1: Comparison of number of faces, $\mathcal{F}$ (for structured scenes) against surface reconstruction approaches. For all scenes, our approach generates a significantly lower number of faces without compromising the scene’s accuracy.
@@
Figure 10: Comparison with surface reconstruction approaches including both structured and non-structured scenes.
@@ 5.4 Comparison with Shape Approximation Approaches
Figure 11: Comparison with shape approximation approaches. The blue region shows the unclipped mesh due to the absence of the ceiling in the original mesh.
As shown in figure 11, all approaches generate structure-preserving mesh. However, focusing on structured scenes, the primary objective of our work, we observe that our approach significantly reduces the number of faces, as detailed in table 3, thus largely simplifying the structured mesh. Unlike other approaches, our method inherently distinguishes between structured and non-structured scenes. Thus, for a fair comparison, table 3 contains a comparison of the number of faces for only the structured scenes. Notably, in scenes such as frl_apartment_0 in figure 11, where ceilings are absent, our approach’s dependence on clipping ceiling planes with respect to neighboring planes may result in unclipped extended planes, as indicated by the blue region.
Additionally, we evaluate RMSE, as shown in table 4, including both structured and non-structured scenes. RMSE values across all approaches are markedly lower compared to surface reconstruction, owing to the inherent nature of shape approximation approaches, which focuses on mesh decimation rather than reconstruction from the original mesh. In simpler scenes like office_4, the lower number of faces in our approach results in lower RMSE. However, as scene complexity increases such as hotel_0, the number of primitives and faces also increases, leading to elevated RMSE values.